OpenStudy (anonymous):

\[\int\limits x(x-1)^5~dx ---> x=1+u\]\[\int\limits (1+u)(u)^5~du\]How to carry on?

OpenStudy (anonymous):

\[\int\limits u ^{5}dx +\int\limits u ^{6}dx\]

OpenStudy (anonymous):

sry, du

OpenStudy (anonymous):

\[\int\limits (u ^{5}+u ^{6})du=\int\limits u ^{5}du+ \int\limits u^{6}du\]

OpenStudy (anonymous):

u^6 /6 + u^7 / 7 +c

OpenStudy (anonymous):

Why u^6?

OpenStudy (anonymous):

u^5 (1+u)= u^6 + u^5

OpenStudy (anonymous):

Oh, Ok.